You are infusing D5W at 125mL/hr. You just hung a new 1 liter bag and the...

  1. You are infusing D5W at 125mL/hr. You just hung a new 1 liter bag and the patient askes how long it will take to complete the bag. You respond _____________hours.

  1. A patient is to receive 1000mL of lactated Ringer’s solution over a 6 hour period. What would you set the rate to on IV pump? __________

  1. You are to administer 500 mL of 0.9% NSS over 6 hours. The drop factor is 20gtt/mL._____________gtt/min.

  1. Administer 250 mL of 0.45% NSS over 5 hours. The drip factor is 60 gtts/ml. _________________________gtt/min

  1. The physician prescribed 1000 mL of RL to infuse over 12 hours. You would give _____________________mL/hour.

  1. To deliver 250 mL of NSS over a 5 hour period you would set the flow rate to deliver _____________mL/hr.

  1. You are to give 500 mL of NSS at 40 mL/hr. The total infusion time would be __________________hours.

  1. Infuse 1000mL of D5W at 125mL/hr. The infusion set delivers 15 gtts/mL. To ensure the solution will infuse over an 8 hour period, how many drops per minutes should the nurse set the rate of the flow? ____________

  1. The physician ordered 500mL of D5W with 30,000 units of heparin to infuse at 10 mL/hr. Calculate the hourly dosage of heparin. ________U/hr in 10 mL.
Answer #1

Ans) 1 litre of D5W @ 125 ml/hr

1000/125 = 8 hours

- 1000 ml RL to be Infused in 6 hrs:

1000/6 = 166.66 ~ 167 ml/hr

- Formula :

Rate = Total volume to be administered × drop factor÷ time

Rate = 500 × 20 ÷ 6 × 60 . (1 hr = 60 mins )

= 10000 /360 = 27.77 ~ 28 gtts/min.

- 250 ml of 0.545 NSS over 5 hrs

250 × 60 /5 × 60 =50 gtts/ min.